# Superconductivity, Anomalous Hall Effect, and Stripe Order in Rhombohedral Hexalayer Graphene

### v1

@@
The emergence of superconductivity from a momentum-space condensed parent state enables Cooper pairing with finite momentum, a phenomenon of great interest due to its direct connection to the superconducting diode effect in the absence of an external magnetic field Lin et al. (2022); Scammell et al. (2021); Yuan and Fu (2021). Given its unique combination of broken symmetries, the SC i phase provides an ideal platform for investigating nonreciprocal superconducting transport at zero magnetic field. While a detailed study of the superconducting diode effect is beyond the scope of this work, it is worth noting that previous reports of a zero-field diode effect also identified the coexistence of superconductivity and a density wave Lin et al. (2022). However, in those cases, the angular symmetry of the density wave could not be resolved due to the constraints imposed by the Hall-bar sample geometry.
@@
Lastly, we highlight the specific correlation between the angular symmetry of superconducting transport and the anisotropy of the metallic phase: the most robust superconducting response consistently occurs along the direction of maximum conductivity (minimum resistivity) observed in the metallic phase above $T_{c}$. While SC ii and SC iii do not exhibit the extreme anisotropy seen in SC i, they share the same angular interplay between superconducting and metallic transport (see Fig. M11). This behavior stands in stark contrast to recent findings in magic-angle twisted trilayer graphene Zhang et al. (2025), where superconductivity is strongest along the direction of minimum conductivity (maximum resistivity) in the normal state. Taken together, these observations establish a fundamental distinction between the superconducting phases in rhombohedral hexalayer graphene and those previously reported in magic-angle graphene moiré systems Cao et al. (2018); Yankowitz et al. (2019); Lu et al. (2019); Park et al. (2021); Hao et al. (2021); Liu et al. (2022).
@@ acknowledgments
J.I.A.L. wishes to acknowledge helpful discussion with Dima Feldman, Oskar Vafek, Aaron Hui, Mathias Scheurer, and Chunli Huang. E.M. and J.I.A.L. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. J.I.A.L. acknowledge partial support from the Air Force Office of Scientific Research. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.

### v2

@@
Such a shared origin could naturally arise from Coulomb-driven instabilities of a valley-polarized Fermi surface, which favor spontaneous condensation in momentum space Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025). Momentum-space condensation has been widely explored in the context of multilayer graphene systems Zhang et al. (2024a); Lin et al. (2023). From a momentum-polarized parent state, the emergence of superconductivity permits finite-momentum Cooper pairing—a scenario directly linked to the superconducting diode effect at zero magnetic field Lin et al. (2022); Scammell et al. (2022); Yuan and Fu (2021). With its distinct combination of broken symmetries, the SC i phase presents an ideal platform to explore nonreciprocal superconducting transport. While a full study of the diode effect lies beyond the scope of this work, we note that prior zero-field diode reports also involved coexisting superconductivity and a density wave Lin et al. (2022). However, in those cases, the angular symmetry of the density wave could not be resolved due to the constraints of Hall-bar geometries.
The existence of valley polarization above the superconducting transition also distinguishes this system from the A-phase of superfluid 3He, where the normal fluid phase preserves time-reversal symmetry.
Taken together, our observations reveal a remarkable and intricate interplay among multiple emergent electronic orders. Above the superconducting transition temperature, the anomalous Hall effect emerges concurrently with pronounced transport anisotropy at the nematic transition $T_{\text{nem}}$, highlighting a metallic phase characterized by spontaneous momentum-space condensation. In the low-temperature regime, magnetic-field–driven hysteretic switching between degenerate superconducting states strongly supports the chiral nature of the superconducting order parameter Han et al. (2024a). The simultaneous onset of superconductivity and stripe charge order naturally leads to a spatial modulation of the superconducting gap, establishing this system as a promising platform for realizing pair density wave states Agterberg et al. (2020). Furthermore, the coexistence of stripe order and the anomalous Hall effect provides compelling experimental evidence for a stripy anomalous Hall crystal Tan and Devakul (2024); Dong et al. (2024a); Soejima et al. (2024); Dong et al. (2024b); Patri et al. (2024); Dong et al. (2024c).
Despite its coexistence and competition with stripe order, the superconducting transition temperature ($T_{c}$) in rhombohedral hexalayer graphene is significantly higher than values reported for other rhombohedral multilayer graphene systems Han et al. (2024a); Zhou et al. (2021b, 2022). This enhanced $T_{c}$ may arise from stronger Coulomb interactions associated with the flatter bands in hexalayer graphene, as suggested by both theoretical predictions and experimental observations  Zhang et al. (2024b); Muten et al. (2021); Pamuk et al. (2017).
Furthermore, previous studies of rhombohedral pentalayer graphene aligned with hexagonal boron nitride (hBN) have also hinted at the presence of an anomalous Hall crystal phase Lu et al. (2024, 2025). Our observation of stripe order coexisting with the anomalous Hall effect within a similar region of the $n$$D$ phase diagram suggests that stripe formation may represent a universal ordering tendency in rhombohedral multilayer graphene under strong displacement fields.
An outstanding signature of the superconducting phase in rhombohedral tetra- and pentalayer graphene is its remarkable stability against large out-of-plane magnetic fields ($B_{\perp}$an et al. (2024a). In rhombohedral hexalayer graphene, however, this stability depends sensitively on the orientation of the applied current relative to the stripe direction. When current flows along the stripe orientation, $R$$T$ curves exhibit a sharp superconducting transition that persists even under a substantial out-of-plane magnetic field of $B_{\perp}=0.3$ T (Fig. 4h). As illustrated in Fig. M12, $R_{\parallel}$ measured along the stripe direction remains nearly zero up to approximately $B_{\perp}=0.75$ T.
In contrast, when current is applied along $\phi=270^{\circ}$—misaligned by approximately $45^{\circ}$ from the stripe direction—the application of $B_{\perp}=-0.3$ T fully suppresses dissipationless superconducting transport at low temperatures (Fig. 4h). Under this field, superconductivity is observed only when the current is perfectly aligned with the stripe orientation. When $B_{\perp}$ is further increased to $-1$ T, superconductivity is completely suppressed, independent of the current direction (Fig. 4i). In this high-field regime, the $R$$T$ curves display an upturn at low temperatures, consistent with the reemergence of a stripe-ordered insulating phase.
@@
Lastly, we comment on the angular interplay between superconductivity and the metallic phase. In SC i, the most robust superconducting response consistently occurs along the direction of maximum conductivity (minimum resistivity) observed in the metallic phase above $T_{c}$. While SC ii and SC iii do not exhibit the extreme anisotropy seen in SC i, they share the same angular interplay between superconducting and metallic transport (see Fig. M13). This behavior stands in stark contrast to recent findings in magic-angle twisted trilayer graphene Zhang et al. (2025), where superconductivity is strongest along the direction of minimum conductivity (maximum resistivity) in the normal state. These observations establish a fundamental distinction between the superconducting phases in rhombohedral hexalayer graphene and those previously reported in magic-angle graphene moiré systems Cao et al. (2018); Yankowitz et al. (2019); Lu et al. (2019); Park et al. (2021); Hao et al. (2021); Liu et al. (2022).
@@ acknowledgments
J.I.A.L. wishes to thank Oskar Vafek for continued support and theoretical guidance in developing the novel scheme of angle-resolved transport measurement. J.I.A.L. also acknowledges helpful discussions with Sankar Das Sarma, Dima Feldman, Aaron Hui, Mathias Scheurer, Chunli Huang, and Boris Shklovskii. E.M. and J.I.A.L. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. J.I.A.L. acknowledge partial support from the Air Force Office of Scientific Research. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.